# Bohr-Sommerfeld type surgeries and Disk Potential

### v1

@@ 2.3. Bohr-Sommerfeld Profile Surgery
We describe a class of Lagrangians which contain a BSP Lagrangian handle. Let $(M,\omega_{M})$ be a compact symplectic manifold and $L$ be an embedded Lagrangian.
@@
A natural question at this stage is - how does the choice of the curve $\gamma$ affects the Hamiltonian isotopy class of a Lagrangian with BSP Handles? We define the notion of relative Hamiltonian isotopy which lets us extend a Hamiltonian isotopy of $\mathcal{L}_{\gamma_{\tau}}$ in $\phi([a,b]\times Z)$ to an isotopy of the closed Lagrangian $L_{\tau}.$
@@
be a Lagrangian isotopy in $M$ which is constant on the boundary $\partial K$. We say that $\Phi$ is a relative Hamiltonian isotopy if there is a family of Hamiltonian functions $H_{\tau}$ such that the isotopy induced from it is $\Phi$ and
@@
From the definition of relative Hamiltonian isotopy, we know there is a family $H_{\tau}$ which realizes the isotopy of $K$ such that $H_{\tau}(\partial K)=0.$ Note that changing $H_{\tau}$ outside a small normal neighborhood of $K_{\tau}$ still preserves the induced isotopy. We can extend Up to changing $H_{\tau}$ outside a small normal neighborhood of $K$, we can extend $H_{\tau}$ by defining it to be 00 on $L$ and extending to the whole of $M$. The Hamiltonian isotopy from this extension would fix $L\setminus U$ and realize the isotopy of $K$ in $U.$
We now explore how the choice of the curve $\gamma$ determines the relative Hamiltonian isotopy class of $\mathcal{L}_{\gamma}$. We use the notion of exact Lagrangian isotopy as in [pol01, § 6.1] to construct Hamiltonian isotopy. Define the following one-form on (,)×$(-\infty,\infty)\times S^{1}$ with coordinates $(s,\theta)$ ,
@@
We will now show that under the hypothesis of this proposition, $\alpha_{\tau}$ is exact for all $\tau$ such that $\alpha_{\tau}=dH_{\tau}$ where $H_{\tau}|_{\partial\mathcal{L}_{\gamma_{\tau}}}\equiv 0.$
We already know that for any curve $c:[0,1]\rightarrow\{r\}\times e^{i\theta}.\Lambda$, we have
@@
[…]
@@
Let $L$ be a Lagrangian with a handle modelled on $\gamma$ with profile $\Lambda\in Z$. Choose a conjugate path $\widetilde{\gamma}$ of the curve $\gamma$. The Bohr-Sommerfeld Profile (BSP) surgery of $L$ with path modelled on $\widetilde{\gamma}$ is defined as
(2.3.2) $BSP^{\Lambda}(L)=(L\setminus\phi^{-1}(L))\cup_{\Lambda_{2}}\mathcal{L}_{\widetilde{\gamma}}$
@@
We call a BSP surgery to be zero-area if the conjugate path $\widetilde{\gamma}$ satisfies
@@
We denote such surgeries as $BSP^{\Lambda}_{0}(L).$
@@
If $L$ is a monotone Lagrangian in $M$, then a zero-area BSP-surgery $BSP_{0}(L)$ is also monotone.
@@
The proof is similar to the proof of [Cha23, Lemma 2.21]. We will show that for any disk $u:\mathbb{D}^{2}\rightarrow(M,BSP_{0}(L))$, there is a disk $u^{\prime}:\mathbb{D}^{2}\rightarrow(M,L)$ such that the Maslov number and symplectic action of $u,u^{\prime}$ are same. Let $[\theta_{1},\theta_{2}]$ be a maximal arc in $\partial\mathbb{D}^{2}$ such that $u([\theta_{1},\theta_{2}])\subset\phi(\mathcal{L}_{\widetilde{\gamma}}).$ We can construct $u^{\prime}$ by constructing a homotopy of paths $\partial u_{[\theta_{1},\theta_{2}]}$ with zero area and zero Maslov index. The existence of such a homotopy follows from
$\int_{\gamma}\lambda_{k}=\int_{\widetilde{\gamma}}\lambda_{k},$
that $Z$ is simply connected and the vanishing of Maslov class on Bohr-Sommerfeld covers (see [rg19, § 4.2]). ∎
2.4. Examples
The primary motivation of behind Bohr-Sommerfeld-Profile surgery is to have a generalized surgery operation on Lagrangians which preserves the Maslov class.
@@ 2.4.1. Polterovich Surgery
Recall the setup for Polterovich surgery ([pol:surg]), let $L_{0}$ be an immersed Lagrangian with a single immersed double point $p$. By a version of Darboux’s theorem one can arrange a chart $(U,\psi)$ around $p$ such that $f^{-1}(L_{0})=\mathbb{R}^{n}\cup i\mathbb{R}^{n}.$ A Polterovich surgery of $L_{0}$ at $p$ replaces $f^{-1}(L_{0})$ with a Lagrangian handle $S^{n-1}\times[0,1]$ and there are two classes, $H_{+},H_{-}$, of such handles.
@@ 2.5. Topology of BSP-Surgeries
Assume that $\Lambda\rightarrow L_{Y}$ is a $k:1$ cover, then the group action $e^{i\frac{2\pi}{k}}$ on $Z$ restricts to a diffeomorphism (deck transformation) of $\Lambda$. Denote this diffeomorphism as $\tau$.
@@
If $\tau$, the restriction of the map $e^{i\frac{2\pi}{k}}:Z\rightarrow Z$ to $\Lambda,$ is homotopic to identity, then the BSP-Surgery preserves the homeomorphism class, i.e.,
$BSP(L)\underbrace{\cong}_{Homeo}L.$
@@
$L\setminus\phi(\Lambda\times[0,1])\bigcup\Lambda\times[0,1]/\sim,$
@@
(2.5.1) $BSP(L)\cong_{\text{Homeo}}L\setminus\phi(\Lambda\times[0,1])\bigcup\Lambda\times[0,1]/\sim,$
@@
From the Lemma 2.5.1 it is enough to show that the restriction of the map $e^{\frac{i2\pi}{k}}$ (which we call $\tau$) to $\Lambda$ is isotopic to identity. Indeed, if it were not the case, then the circle bundle $\Lambda\times[0,1]/\sim$ where the equivalence relation $\sim$ is $(p,0)\sim(\tau(p),1)$ over $L_{Y}$ will be a non-trivial circle bundle. We know that the restriction $Z|_{L_{Y}}$ is a trivial bundle. Indeed, since $H^{2}(L_{Y})$ is torsion-free and the Euler class of $Z|_{L_{Y}}$ is given by the curvature form $\omega_{Y}$ which vanishes on $L_{Y}$, we have that $[\omega_{Y}]=0$ in $H^{2}(L_{Y},\mathbb{Z})$. This contradiction implies that $\tau$ is homotopic to identity and thus
@@
Note that if we take $L_{Y}$ to be an even-dimensional real projective space, $\mathbb{R}P^{2n}$ as a monotone submanifold in $\mathbb{C}P^{2n}$, then the Bohr-Sommerfeld surgery with profile the real Legendrian sphere $S^{2n}$ in $S^{4n+1}$ changes the topological type. This follows from the fact that the map $\tau$ in Lemma 2.5.1 is the antipodal map, which is not isotopic to identity. This is a shadow of a similar phenomenon in Polterovich’s surgery : Polterovich-surgeries of two odd-dimensional Lagrangians $L_{1},L_{2}$ admit two distinct resolutions, $L_{1}\#L_{2}\#P$ and $L_{1}\#L_{2}\#Q$. Here
@@ 3. Zero Area BSP and Disk Potential
As we have seen in Section 2, Bohr-Sommerfeld-Profile surgeries can construct homeomorphic Lagrangians. One of the main motivations behind the construction of Bohr-Sommerfeld-Profile (BSP) surgery is to construct newer symplectic classes of Lagrangians. We will study the effect of BSP surgery on disk-potential, a powerful invariant of oriented compact spin monotone Lagrangians. The disk-potential has been used to distinguish between the infinitely many Vianna tori [Via16, Via17] and their lifts [chw, DTVW]. We provide a change of potential formula for BSP-surgery under certain assumptions, see Assumptions 3.5.
@@ 3.1. Background on disk-potential
Assume $L_{Y}$ is a compact, oriented, spin monotone Lagrangian in $Y$. Choose $J$ to be a compatible almost complex structure on $Y$. Let $\mathcal{M}(\beta,J)$ denote the moduli space of $J$-holomorphic disks with boundary on $L_{Y}$ in class $\beta\in H_{2}(Y,L_{Y})$ with one boundary marking. Denote the Maslov functional on $H_{2}(Y,L_{Y})$ as $\mu$.
There is a natural evaluation map, $ev_{\beta}$, associated with $\mathcal{M}(\beta,J)$. The disk potential $W_{L_{Y}}$ is a polynomial obtained from counting Maslov 2 $J$-holomorphic disks passing through a generic point on $L_{Y}$. Choose a basis $(e_{1},\dots,e_{k})$ of $H_{1}(L_{Y})$.
$W_{L_{Y}}:\text{Hom}(H_{1}(L_{Y}),\mathbb{C}^{*})$ $\longrightarrow\mathbb{C}$
$W_{L_{Y}}(x_{1},\dots,x_{k})$ $=\sum_{\beta\in H_{2}(Y,L_{Y}),\mu(beta)=2}(sgn)n_{\beta}x^{\partial\beta},$
@@
(3.1.1) $x^{-v}W_{L_{Y}}:\text{Hom}(H_{1}(L_{Y}),\mathbb{C}^{*})\rightarrow\mathbb{C}$
has a lift
(3.1.2) $W_{\Lambda}:\text{Hom}(H_{1}(\Lambda),\mathbb{C}^{*})\rightarrow\mathbb{C}.$
This follows from the fact that if $v^{\prime}$ is another cycle bounding a Maslov two disk, then the cycle $v^{\prime}-v$ has a horizontal lift to a cycle in $\Lambda$. We call $W_{\Lambda}$ the augmentation polynomial of the lifted Legendrian $\Lambda$. See [bcsw3, §2.1] for results relating the augmentation polynomial with the Chekanov-Eliashberg algebra, $CE(\Lambda),$ of the Legendrian $\Lambda.$
@@
$\left(\frac{1}{x_{1}\dots x_{n}}\right)^{-1}W_{Cl}=(1+x_{1}^{2}x_{2}\dots x_{n}+x_{1}x_{2}^{2}\dots x_{n}+\dots)$
$\text{setting }y_{k}\text{ such that }p_{*}(y_{k})=(x_{1}..x_{k}^{2}..x_{n}),\;\text{ we have}$
$W_{\Lambda_{Cl}}=(1+y_{1}+\dots+y_{n})$
@@
We say an almost complex structure $J$ on $Y$ to be regular for disk-counting in $L_{Y}$ if the moduli space of $J-$holomorphic disks with boundary on $L_{Y}$ is regularly cut out for all classes of Maslov index less or equal to two.
Definition 3.4 (Enough Regular Disks (ERD) Property).
We say a monotone compact oriented spin Lagrangian $L_{Y}$ has the enough regular disks (ERD) property if there is a choice of almost complex structure $J$ that regular for disk-counting such that there for any $p\in L_{Y}$ and any $m>0$ there is a regular $J-$holomorphic disk $u$ of Maslov number $2m$ passing through the point $p.$
@@
We make some simplifying assumptions on the Lagrangian $L$ which helps us in comparing disk-potentials:
@@
  6. A.6
    The Lagrangian $L_{Y}\hookrightarrow Y$ has enough regular disk (ERD) property.
For the remainder of the section we will suppose that the Lagrangian $L$, on which we perform Zero-BSP surgery, satisfies Assumption 3.5.
Notice that if $L$ satisfies all the assumptions in Assumptions 3.5, the zero-area BSP surgery, $BSP_{0}(L)$ also satisfies them.

### v2

@@ 2.3. Bohr-Sommerfeld Profile Surgery
As we have already mentioned, BSP surgery is obtained by switching exact fillings. We now describe a class of Lagrangians on which we can perform a BSP surgery. These Lagrangians are obtained by gluing Lagrangian fillings and cappings of the Legendrian $\Lambda_{2}$.
Let $(M,\omega_{M})$ be a compact symplectic manifold and $L$ be an embedded Lagrangian.
@@
A natural question at this stage is, how does the choice of the curve $\gamma$ affect the Hamiltonian isotopy class of a Lagrangian-with-BSP handle? We now investigate this line of thought. We define the notion of relative Hamiltonian isotopy which lets us extend a Hamiltonian isotopy of $\mathcal{L}_{\gamma_{\tau}}$ in $\phi([a,b]\times Z)$ to an isotopy of the closed Lagrangian $L_{\tau}.$
@@
be a Lagrangian isotopy of a Lagrangian $K$ in $M$ which is constant on the boundary $\partial K$. We say that $\Phi$ is a relative Hamiltonian isotopy if there is a family of Hamiltonian functions $\{H_{\tau}\}_{\tau\in[0,1]}$ such that the isotopy induced from it is $\Phi$ and $H_{\tau}$ is equal to 0 on the boundary $\partial K$, i.e.
@@
From the definition of relative Hamiltonian isotopy, we know there is a family $H_{\tau}$ which realizes the isotopy of $K$ such that $H_{\tau}(\partial K)=0.$ Note that changing $H_{\tau}$ outside a small normal neighborhood of $K_{\tau}$ still preserves the induced isotopy. We can extend upto changing $H_{\tau}$ outside a small normal neighborhood of $K$, we can extend $H_{\tau}$ by defining it to be 0 on $L$ and extending to the whole of $M$. The Hamiltonian isotopy from this extension would fix $L\setminus U$ and realize the isotopy of $K$ in $U.$
We now explore how the choice of the curve $\gamma$ determines the relative Hamiltonian isotopy class of $\mathcal{L}_{\gamma}$. We use the notion of exact Lagrangian isotopy as in [pol01, § 6.1] to construct Hamiltonian isotopy. Define the following one-form on $(-\infty,\infty)\times S^{1}$ with coordinates $(s,\theta)$ ,
@@
We will now show that under the hypothesis of this proposition, $\alpha_{\tau}$ is exact for all $\tau$ such that $\alpha_{\tau}=dH_{\tau}$ where $H_{\tau}|_{\partial\mathcal{L}_{\gamma_{\tau}}}\equiv 0.$ We already know that for any curve $c:[0,1]\rightarrow\{r\}\times e^{i\theta}.\Lambda$, we have
@@
$\int_{P_{\tau}}\alpha_{\tau}=0\iff\frac{d}{d\tau}\int_{0}^{1}e^{\gamma^{\tau}_{l}(t)/k}{{\gamma^{\tau}}^{\prime}}_{\theta}(t)dt=0\iff\int_{\gamma_{\tau}}\lambda_{k}\text{ is constant}$
@@
Let $L$ be a Lagrangian with a handle modelled on $\gamma$ with profile $\Lambda\subset Z$. Choose a conjugate path $\widetilde{\gamma}$ of the curve $\gamma$. The Bohr-Sommerfeld Profile (BSP) surgery of $L$ with path modelled on $\widetilde{\gamma}$ is defined as
(2.3.2) $BSP^{\Lambda}(L)=(L\setminus\phi^{-1}(L))\cup_{\phi(\partial\mathcal{L}_{\gamma})}\mathcal{L}_{\widetilde{\gamma}}$
@@
We call a BSP surgery to be of zero-area if the conjugate path $\widetilde{\gamma}$ satisfies
@@
We denote the result of such a surgery as $BSP^{\Lambda}_{0}(L)\subset M.$
Definition 2.22.
[…]
Remark 2.23.
Note that Reeb fillability is equivalent to the following condition: for every Reeb chord $c$ with endpoints in $\Lambda_{2}$, there is a smooth punctured disk, i.e. a map from $\mathbb{D}^{2}\setminus\{1\}$, in $\mathbb{R}\times Z$ with boundary on $\mathcal{L}_{\gamma}$ which is asymptotic to $c$ near the boundary puncture.
@@
If $L$ is a monotone Lagrangian in $M$, then a zero-area BSP-surgery with a Reeb fillable Legendrian profile $BSP_{0}(L)$ is also monotone.
@@
The proof is similar to the proof of [Cha23, Lemma 2.21]. We will show that for any disk $u:\mathbb{D}^{2}\rightarrow(M,BSP_{0}(L))$, there is a disk $u^{\prime}:\mathbb{D}^{2}\rightarrow(M,L)$ such that the Maslov number and symplectic action of $u,u^{\prime}$ are same. This will be enough for the proof since we know $L$ is monotone.
[…]
We will construct $u^{\prime}$ by redefining $u$ on $U_{[\theta_{1},\theta_{2}]}$ such that $u^{\prime}=u$ on $\alpha_{[\theta_{1},\theta_{2}]}$. Note that since Bohr-Sommerfeld covers are Maslov 0 from [rg19, § 4.2], it is enough to check that such a $u^{\prime}$ can be constructed. The equality of symplectic area of $u$ and $u^{\prime}$ would then follow from the exactness of $\mathcal{L}_{\gamma},\mathcal{L}_{\widetilde{\gamma}}$ and
$\int_{\gamma}\lambda_{k}=\int_{\widetilde{\gamma}}\lambda_{k}.$
Since we assumed $\Lambda$ was Reeb fillable, we see that such a $u^{\prime}$ exists.
2.4. Examples and relations to other surgeries
The primary motivation behind introducing Bohr-Sommerfeld-Profile surgery is to have a generalized surgery operation on Lagrangians which preserves the Maslov class. We now recover classical Polterovich surgery on Lagrangians from Bohr-Sommerfeld profile surgeries.
@@ 2.4.1. Polterovich Surgery
The setup for Polterovich surgery ([pol:surg]) starts off with an immersed Lagrangian $L_{0}$ that has a single immersed double point $p$. By a version of Darboux’s theorem one can arrange a chart $(U,\psi)$ around $p$ such that $f^{-1}(L_{0})=\mathbb{R}^{n}\cup i\mathbb{R}^{n}.$ Recall that a Polterovich surgery of $L_{0}$ at $p$ replaces $f^{-1}(L_{0})$ with a Lagrangian handle $S^{n-1}\times[0,1]$ and there are two classes, $H_{+},H_{-}$, of such handles.
@@ 2.5. Topology of BSP-Surgeries
Assume that the Bohr-Sommerfeld Legendrian $\Lambda$ is a $k$-fold cover of $L_{Y}$ via the restriction of the projection map $Z\rightarrow Y$. Then, the action of the element $e^{i\frac{2\pi}{k}}\in S^{1}$ on $Z$ restricts to a diffeomorphism (and a deck-transformation over $L_{Y}$) of $\Lambda$. We denote this diffeomorphism as $\tau$.
@@
If $\tau:\Lambda\rightarrow\Lambda$, the multiplication map defined by $p\mapsto e^{i\frac{2\pi}{k}}\cdot p$ to $\Lambda,$ is homotopic to the identity map, then the BSP-Surgery preserves the homeomorphism class, i.e.,
$BSP(L)\cong_{homeo}L.$
@@
$L\setminus\phi(\Lambda\times[0,1])\bigcup(\Lambda\times[0,1])/\sim,$
@@
(2.5.1) $BSP(L)\cong_{\text{Homeo}}L\setminus\phi(\Lambda\times[0,1])\bigcup(\Lambda\times[0,1])/\sim,$
@@
From Lemma 2.5.1 it is enough to show that the restriction of the multiplication by $e^{\frac{i2\pi}{k}}$ map (which we call $\tau$) to $\Lambda$ is isotopic to identity. Indeed, if it were not the case, then the circle bundle $\Lambda\times[0,1]/\sim$ where the equivalence relation $\sim$ is $(p,0)\sim(\tau(p),1)$ over $L_{Y}$ will be a non-trivial circle bundle. We know that the restriction $Z|_{L_{Y}}$ is a trivial bundle. Indeed, since $H^{2}(L_{Y})$ is torsion-free and the Euler class of $Z|_{L_{Y}}$ is given by the curvature form $\omega_{Y}$ which vanishes on $L_{Y}$, we have that $\omega_{Y}]=0$ in $H^{2}(L_{Y},\mathbb{Z})$. This contradiction implies that $\tau$ is homotopic to identity and thus
@@
Note that if we take $L_{Y}$ to be an even-dimensional real projective space, $\mathbb{R}P^{2n}$ as a monotone submanifold in $\mathbb{C}P^{2n}$, then the Bohr-Sommerfeld surgery with profile the real Legendrian sphere $S^{2n}$ in $S^{4n+1}$ changes the topological type. This follows from the fact that the map $\tau$ in Lemma 2.5.1 is the antipodal map, which is not isotopic to identity. This is a shadow of a similar phenomenon in Polterovich’s surgery : Polterovich-surgeries of two odd-dimensional Lagrangians $L_{1},L_{2}$ admit two distinct smoothings, $L_{1}\#L_{2}\#P$ and $L_{1}\#L_{2}\#Q$. Here
@@ 3. Zero Area BSP and Disk Potential
As we have seen in Section 2.5, Bohr-Sommerfeld-Profile surgeries can construct homeomorphic Lagrangians. One of the main application of Bohr-Sommerfeld-Profile (BSP) surgery is to construct exotic symplectic classes of Lagrangians. We will study the effect of BSP surgeries on the disk-potential of oriented compact spin monotone Lagrangians. The disk-potential has been used to distinguish between the infinitely many Vianna tori [Via16, Via17] and their lifts [chw, DTVW]. We provide a change of potential formula for BSP-surgery under certain assumptions, see Assumptions 3.4.
@@ 3.1. Background on disk-potential
Assume $L_{Y}$ is a compact, oriented, spin monotone Lagrangian in $Y$. Choose $J$ to be a compatible almost complex structure on $Y$. We denote the moduli space of $J$-holomorphic disks with boundary on $L_{Y}$ in class $\beta\in H_{2}(Y,L_{Y})$ with one boundary marking by $\mathcal{M}(\beta,J)$. Let $\mu$ be the Maslov functional on $H_{2}(Y,L_{Y})$.
For a generic choice of $J$, the moduli space of once boundary marked $J$-holomorphic disks of Maslov number 2 is a compact oriented manifold. There is a natural evaluation map, $ev_{\beta}$, associated with $\mathcal{M}(\beta,J)$ which is defined by evaluation at the boundary marking. The disk potential $W_{L_{Y}}$ is a polynomial obtained from counting Maslov 2 $J$-holomorphic disks passing through a generic point on $L_{Y}$. Choose a basis $(e_{1},\dots,e_{k})$ of $H_{1}(L_{Y})$.
$W_{L_{Y}}:$ $\mathrm{Hom}\big(H_{1}(L_{Y}),\mathbb{C}^{*}\big)\longrightarrow\mathbb{C},$
$(x_{1},\dots,x_{k})\longmapsto\sum_{\begin{subarray}{c}\beta\in H_{2}(Y,L_{Y})\\ \mu(\beta)=2\end{subarray}}(\mathrm{sgn})\,n_{\beta}\,x^{\partial\beta}.$
@@
(3.1.1) $x^{-v}W_{L_{Y}}:\mathrm{Hom}\big(H_{1}(L_{Y}),\mathbb{C}^{*}\big)\longrightarrow\mathbb{C}$
lifts to
(3.1.2) $W_{\Lambda}:\mathrm{Hom}\big(H_{1}(\Lambda),\mathbb{C}^{*}\big)\longrightarrow\mathbb{C}.$
This follows from the fact that if $v^{\prime}$ is another cycle bounding a Maslov two disk, then the cycle $v^{\prime}-v$ has a horizontal lift [^4] Here we are using the monotonicity of our set up. to a cycle in $\Lambda$. We call $W_{\Lambda}$ the augmentation polynomial of the lifted Legendrian $\Lambda$. See [bcsw3, §2.1] for results relating the augmentation polynomial with the Chekanov-Eliashberg algebra, $CE(\Lambda),$ of the Legendrian $\Lambda.$
@@
$\left(\frac{1}{x_{1}\cdots x_{n}}\right)^{-1}W_{Cl}$ $=1+x_{1}^{2}x_{2}\cdots x_{n}+x_{1}x_{2}^{2}\cdots x_{n}+\dots,$
$\text{set }y_{k}\text{ such that }\quad p_{*}(y_{k})$ $=x_{1}\cdots x_{k}^{2}\cdots x_{n},$
$\text{then}\quad W_{\Lambda_{Cl}}$ $=1+y_{1}+\dots+y_{n}.$
Here $y_{k}$ are cycles in $T_{Cl}$ that lift to a basis of the Legendrian tori.
@@
We call an almost complex structure $J$ on $Y$ to be regular for disk-counting in $L_{Y}$ if the moduli space of $J$-holomorphic disks with boundary on $L_{Y}$ is regularly cut out for all classes of Maslov index less or equal to two.
@@
We state the assumptions on the Lagrangian $L$ which are essential for our proof strategy :
@@
    $\Lambda$ is a Bohr-Sommerfeld lift of a Lagrangian $L_{Y}$ that has minimal Maslov number 2.
  3. A.3
@@
Remark 3.5.
One can recover a similar disk potential relation by relaxing A.6 above which allows the for the homeomorphism class of the Lagrangian to change while preserving the splitting property in A.3. We do not have applications of such a result in our article so we avoid pursuing such a result.
For the remainder of the section we will suppose that the Lagrangian $L$, on which we perform Zero-BSP surgery, satisfies Assumption 3.4. Notice that if $L$ satisfies all the assumptions in Assumptions 3.4, the zero-area BSP surgery, $BSP_{0}(L)$ also satisfies them.